# Inverse Hamiltonian reduction in type A and generalized slices in the affine Grassmannian

### v1

@@ 2.4. Matrix realization of generalized slices
$\operatorname{Bun}_{B}(\mathbb{P}^{1})\xrightarrow{\cong}\left[B{{[\![t^{-1}]\!]}}\backslash B{{(\!(t^{-1})\!)}}/B[t]\right]\qquad\text{and}\qquad\operatorname{Bun}_{G}(\mathbb{P}^{1})_{(B,\infty)}\xrightarrow{\cong}\left[I_{\infty}\backslash G{{(\!(t^{-1})\!)}}/G[t]\right]$
@@
$\operatorname{Bun}_{B}(\mathbb{P}^{1})\to\operatorname{Bun}_{G}(\mathbb{P}^{1})_{(B,\infty)}\quad\quad\text{is modeled by}\quad\quad\left[B{{[\![t^{-1}]\!]}}\backslash B{{(\!(t^{-1})\!)}}/B[t]\right]\to\left[I_{\infty}\backslash G{{(\!(t^{-1})\!)}}/G[t]\right]\ .$
@@ 2.5. Open Zastava and weakly antidominant slices
[…]
@@ 2.6. Coulomb branch description of generalized slices
¯$\underline{\mathcal{M}}_{C}({\bf{G}},{\bf{N}})=\operatorname{Spec}H_{\bullet}^{{\bf G}(\mathcal{O})\times{\bf G}_{F}(\mathcal{O})}(\mathcal{R}({\bf{G}},{\bf{N}}))\qquad\text{and}\qquad\underline{\mathcal{M}}_{C}({\bf{G}},{\bf{N}})_{\mathfrak{t}_{F}}=\operatorname{Spec}H_{\bullet}^{{\bf G}(\mathcal{O})\times{\bf T}_{F}(\mathcal{O})}(\mathcal{R}({\bf{G}},{\bf{N}}))$
@@
¯$\underline{\mathcal{M}}_{C}({\bf{G}},{\bf{N}})_{\mathfrak{t}_{F}}\xrightarrow{\cong}\underline{\textup{Gr}}^{\underline{\lambda}}_{\mu}\ .$
@@
¯$\underline{\mathcal{M}}_{C}({\bf{G}},{\bf{N}})\xrightarrow{\cong}\underline{\textup{Gr}}_{\mu,\mathbb{A}^{N}/\mathfrak{S}_{\underline{\lambda}}}^{\underline{\lambda}}\ .$
@@
3. Applications to Lie theory in type $A$
Let $\mu=(\mu_{1}\geq...\geq\mu_{n-1}\geq 0)$ be a dominant coweight of ${{\textup{PGL}}}_{n}$, the set of which we identify with partitions of length at most $n-1$, where each such $\mu$ is a partition of the integer $N=|\mu_{i}|=\sum_{i\in I}\mu_{i}$.
@@ 3.1. Nilpotent orbits and their transverse slices
Let ${\mathcal{N}}\subset\mathfrak{g}^{*}$ be the nilcone of $\mathfrak{g}^{*}$. The nilpotent coadjoint orbits in $\mathfrak{g}^{*}$ are in bijection with the partitions of $n-1$ (according to the Jordan type). Let $\mu$ be such a partition, as above, and let $\chi_{\mu}=\langle f_{\mu},\cdot\rangle$ be an element of the nilpotent coadjoint orbit $\mathbb{O}_{\mu}$ determined by $f_{\mu}\in{{\mathfrak{g}}}$ a nilpotent of Jordan type $\mu$. Note that the dominance order on coweights induces the order on partitions corresponding to the closure order on orbits.
@@
(3.1.5) $t\cdot s=t^{2-\rm{ad}_{h}}s$
Note that the action, contracts to the fixed point $\chi\in S_{\mu}$. This action of $\mathbb{G}_{m}$ scales the Poisson bracket with weight $-1$.
@@ 3.2. The Mirković–Vybornov isomorphism
Let $\mu=(\mu_{1}\geq...\geq\mu_{n-1}\geq 0)$ be a dominant coweight of ${{\textup{PGL}}}_{n}$, viewed as a partition of the integer $N=|\mu_{i}|=\sum_{i\in I}\mu_{i}$ of length at most $n-1$, as above. Let $\lambda=N\omega_{1}\in\Lambda_{{{\textup{PGL}}}_{n}}^{+}$ so that following the notation of Equation 2.6.1 we have
@@
so that $p_{i}:=m_{i}-m_{i-1}=\mu_{n-i}$ for $i=2,...,n-1$, $p_{1}:=m_{1}=0$, $p_{n}:=N-m_{n-1}=\mu_{1}$, and in summary $\pi:=(p_{1}\leq p_{2}\leq\cdots\leq p_{n})$ the partition of Section 2.1 in [WWY20] is simply the partition $\mu$ thought of in ascending order.
@@
Note that in writing $\mu=(\mu_{1}\geq...\geq\mu_{n-1})$ we have departed from the conventions of loc. cit.; their convention for the definition of $\mu_{i}$ is the one which we follow for $\lambda$ in writing $\lambda=\sum_{i=1}^{n-1}\lambda_{i}\omega_{n-i}$. Our notation has the benefit that $\mu_{i}$ simply give the components of the partition $\pi$, as noted above, which labels the relevant nilpotent orbit in the following theorem. Thus, we do not refer to the latter notation, and since we essentially always take $\lambda=N\omega_{1}$ or 00 in the following applications, this discrepancy in convention between $\lambda$ and $\mu$ will not be especially relevant.
@@
¯$\underline{\mathcal{M}}_{C}({\bf{G}},{\bf{N}})_{\mathfrak{t}_{F}}\xrightarrow{\cong}\underline{\textup{Gr}}_{\mu}^{{\underline{\lambda}}_{N}}\xrightarrow{\cong}S_{\mu}\times_{\mathfrak{h}/W}\mathfrak{h}\qquad\text{and}\qquad\underline{\mathcal{M}}_{C}({\bf{G}},{\bf{N}})\xrightarrow{\cong}\underline{\textup{Gr}}_{\mu,\mathbb{A}^{N}/\mathfrak{S}_{N}}^{{\underline{\lambda}}_{N}}\xrightarrow{\cong}S_{\mu}\ ,$
@@
[…]
@@ 3.3. Equivariant Slodowy slices
3.4. Zastava presentation of equivariant Slodowy slices in type $A$
The cotangent bundle $T^{*}G$ for $G={{\textup{GL}}}_{N}$ has a realization as a Zastava space $Z^{\alpha_{0}^{*}}_{{{\textup{PGL}}}_{2N}}$ by [BP08], thus in particular a generalized slice in the affine Grassmannian of ${{\textup{PGL}}}_{2N}$. Further, by the Theorem 2.15 recalled above from [BFN19a], $T^{*}G$ is also realized as the Coulomb branch of a type $A_{2N}$ quiver, a linear quiver of length $12N-$ 1. We now state this formally for future reference:
@@ 3.4. Zastava presentation of equivariant Slodowy slices in type $A$
where ${\bf{G}}$ and ${\bf{N}}$ are as in the preceding Theorem 3.6, and in turn the definitions recalled in briefly in Section 2.6 above from [BFN18]. Since the resulting representation is trivial on the natural diagonal character ×$\mathbb{C}^{\times}\to{\bf{G}}={{\textup{GL}}}(V)=\prod_{i=1}^{N}{{\textup{GL}}}_{i}$, the above construction naturally descends to define a $\widetilde{\bf{G}}$-equivariant map $\mathcal{R}(\widetilde{\bf{G}},{\bf{N}})\xrightarrow{\tilde{\pi}_{F}}\mathrm{Gr}_{{{\textup{PGL}}}_{N}}$ where $\widetilde{\bf{G}}={\bf{G}}/\mathbb{C}^{\times}$. In summary, we obtain sheaves
@@
[…]
@@
To begin, note that there is a manifest ${{\textup{GL}}}_{N}$ action on $\mathcal{M}_{C}({\bf{G}},{\bf{N}})$ as a factor of the group $L_{\textup{bal}}$ corresponding to the balanced subquiver $(N,N-1,N-2,...,1)$. Moreover, this defines a trivial ${{\textup{GL}}}_{N}$-torsor, as we can directly calculate
@@
We also recall the following definition of the general genus zero Moore–Tachikawa varieties, with punctures labeled by dominant coweights $\mu^{1},...,\mu^{b}\in\Lambda^{+}_{{{\textup{PGL}}}_{N}}$ with $|\mu^{k}|=N$ for each $k=1,...,b$, or $\mu^{k}=0$ in which case the puncture is called ‘full’ and corresponds to those discussed in [BFN19b]:
@@ 4.1. Quantization in the algebraic setting
First, let us review some preliminaries on quantization. We largely follow the language of [BK04a, Los12]. For now, suppose that $Y$ is a smooth, symplectic, affine scheme over $\mathbb{\mathbb{C}}$. By definition, ${\mathcal{O}}_{Y}$ is a sheaf of Poisson algebras over $Y$. Suppose also, that $Y$ has an action of $\mathbb{G}_{m}$ such that the Poisson tensor is scaled with weight [^1] Usually one asks that $\mathbb{G}_{\hbar}$ scales the symplectic form and $\hbar$ with weight $2$, but one is then forced to change variables to $\sqrt{\hbar}$. We opt to avoid this by halving the weights. In characteristic zero, this amounts to little more than a matter of convention. $-1$, we use $\mathbb{G}_{\hbar}$ for this distinguished action. This group also also acts on $\mathbb{\mathbb{C}}{{[\![\hbar]\!]}}$ by scaling $\hbar$ with unit weight. Much of this discussion holds in a more general setting but in the spirit of brevity, we specialize to the setting we need.
A formal quantization of the Poisson algebra ${\mathcal{O}}(Y)$ is a pair $(A,\varphi)$, where $A$ is an associative $\mathbb{\mathbb{C}}{{[\![\hbar]\!]}}$ algebra, flat over $\mathbb{\mathbb{C}}{{[\![\hbar]\!]}}$ and complete in the $\hbar$-adic topology and $\varphi:A/\hbar A\rightarrow{\mathcal{O}}(Y)$ is an isomorphism of Poisson algebras.
In particular, we shall be interested in graded quantizations. Moreover, this group acts on $\mathbb{\mathbb{C}}{{[\![\hbar]\!]}}$ by scaling $\hbar$ with weight $1$. A formal quantization $(A,\varphi)$ is graded if $\mathbb{G}_{\hbar}$ acts by algebra automorphisms and $\varphi:A/\hbar A\rightarrow{\mathcal{O}}(Y)$ is $\mathbb{G}_{\hbar}$-equivariant.
Often, we will omit the isomorphism $\varphi$ in our notation when it is not relevant.
These definitions have natural sheaf-theoretic analogues in quantizations of the structure sheaf ${\mathcal{O}}_{Y}$.
A formal quantization of ${\mathcal{O}}_{Y}$ is a sheaf of associative, flat, $\mathbb{\mathbb{C}}{{[\![\hbar]\!]}}$ algebras, ${\mathcal{A}}$, on $Y$ that is complete in the $\hbar$-adic topology and with an isomorphism ${\mathcal{A}}/\hbar{\mathcal{A}}\rightarrow{\mathcal{O}}_{Y}$ of sheaves of Poisson algebras.
A formal quantization of ${\mathcal{O}}_{Y}$ is graded if $Y$ has an action of $\mathbb{G}_{\hbar}$ and ${\mathcal{A}}$ is a $\mathbb{G}_{\hbar}$-equivariant sheaf and $\varphi$ is $\mathbb{G}_{\hbar}$-equivariant. Note that the global sections of a graded quantization of ${\mathcal{O}}_{Y}$ are a graded quantization of ${\mathcal{O}}(Y)$.
@@
The moduli space of isomorphism classes of graded quantizations of ${\mathcal{O}}_{Y}$ are isomorphic to the second de Rham cohomology $H^{2}_{\rm dR}(Y)$
This is the familiar result for quantizations of symplectic manifolds in the $C^{\infty}$ setting, see for example [dWL83, Fed94]. The result holds, in the algebraic context, for more general schemes but we only need the affine case.

### v2

@@ 2.4. Matrix realization of generalized slices
[…]
@@
[…]
@@ 2.5. Open Zastava and weakly antidominant slices
$B=\mathcal{P}_{B}|_{\{\infty\}}\to(\mathcal{P}_{B}\times_{B}G)|_{\{\infty\}}\xrightarrow{\varphi}G$
@@ 2.6. Coulomb branch description of generalized slices
$\underline{\mathcal{M}}_{C}({\bf{G}},{\bf{N}})=\operatorname{Spec}H_{\bullet}^{{\bf G}(\mathcal{O})\times{\bf G}_{F}(\mathcal{O})}(\mathcal{R}({\bf{G}},{\bf{N}}))\qquad\text{and}\qquad\underline{\mathcal{M}}_{C}({\bf{G}},{\bf{N}})_{\mathfrak{t}_{F}}=\operatorname{Spec}H_{\bullet}^{{\bf G}(\mathcal{O})\times{\bf T}_{F}(\mathcal{O})}(\mathcal{R}({\bf{G}},{\bf{N}}))$
@@
$\underline{\mathcal{M}}_{C}({\bf{G}},{\bf{N}})_{\mathfrak{t}_{F}}\xrightarrow{\cong}\underline{\textup{Gr}}^{\underline{\lambda}}_{\mu}\ .$
@@
$\underline{\mathcal{M}}_{C}({\bf{G}},{\bf{N}})\xrightarrow{\cong}\underline{\textup{Gr}}_{\mu,\mathbb{A}^{N}/\mathfrak{S}_{\underline{\lambda}}}^{\underline{\lambda}}\ .$
@@
3. Applications to Lie theory in type AA
Let $\mu=(\mu_{1}\geq...\geq\mu_{n-1}\geq 0)$ be a dominant coweight of ${{\textup{PGL}}}_{n}$, the set of which we identify with partitions of length at most n1n-1, where each such $\mu$ is a partition of the integer $N=|\mu_{i}|=\sum_{i\in I}\mu_{i}$.
@@ 3.1. Nilpotent orbits and their transverse slices
Let ${\mathcal{N}}\subset\mathfrak{g}^{*}$ be the nilcone of $\mathfrak{g}^{*}$. The nilpotent coadjoint orbits in $\mathfrak{g}^{*}$ are in bijection with the partitions of n1n-1 (according to the Jordan type). Let $\mu$ be such a partition, as above, and let $\chi_{\mu}=\langle f_{\mu},\cdot\rangle$ be an element of the nilpotent coadjoint orbit $\mathbb{O}_{\mu}$ determined by $f_{\mu}\in{{\mathfrak{g}}}$ a nilpotent of Jordan type $\mu$. Note that the dominance order on coweights induces the order on partitions corresponding to the closure order on orbits.
@@
(3.1.5) $t\cdot s=t^{1-\tfrac{1}{2}\rm{ad}_{h}}s$
Note that the action, contracts to the fixed point $\chi\in S_{\mu}$. This action of $\mathbb{G}_{m}$ scales the Poisson bracket with weight 1-1.
@@ 3.2. The Mirković–Vybornov isomorphism
Let $\mu=(\mu_{1}\geq...\geq\mu_{n-1}\geq 0)$ be a dominant coweight of ${{\textup{PGL}}}_{n}$, viewed as a partition of the integer $N=|\mu_{i}|=\sum_{i\in I}\mu_{i}$ of length at most n1n-1, as above. Let $\lambda=N\omega_{1}\in\Lambda_{{{\textup{PGL}}}_{n}}^{+}$ so that following the notation of Equation 2.6.1 we have
@@
so that $p_{i}:=m_{i}-m_{i-1}=\mu_{n-i}$ for i=2,,n1i=2,...,n-1, $p_{1}:=m_{1}=0$, $p_{n}:=N-m_{n-1}=\mu_{1}$, and in summary $\pi:=(p_{1}\leq p_{2}\leq\cdots\leq p_{n})$ the partition of Section 2.1 in [WWY20] is simply the partition $\mu$ thought of in ascending order.
@@
Note that in writing $\mu=(\mu_{1}\geq...\geq\mu_{n-1})$ we have departed from the conventions of loc. cit.; their convention for the definition of $\mu_{i}$ is the one which we follow for $\lambda$ in writing $\lambda=\sum_{i=1}^{n-1}\lambda_{i}\omega_{n-i}$. Our notation has the benefit that $\mu_{i}$ simply give the components of the partition $\pi$, as noted above, which labels the relevant nilpotent orbit in the following theorem. Thus, we do not refer to the latter notation, and since we essentially always take $\lambda=N\omega_{1}$ or 0 in the following applications, this discrepancy in convention between $\lambda$ and $\mu$ will not be especially relevant.
@@
$\underline{\mathcal{M}}_{C}({\bf{G}},{\bf{N}})_{\mathfrak{t}_{F}}\xrightarrow{\cong}\underline{\textup{Gr}}_{\mu}^{{\underline{\lambda}}_{N}}\xrightarrow{\cong}S_{\mu}\times_{\mathfrak{h}/W}\mathfrak{h}\qquad\text{and}\qquad\underline{\mathcal{M}}_{C}({\bf{G}},{\bf{N}})\xrightarrow{\cong}\underline{\textup{Gr}}_{\mu,\mathbb{A}^{N}/\mathfrak{S}_{N}}^{{\underline{\lambda}}_{N}}\xrightarrow{\cong}S_{\mu}\ ,$
@@
(3.2.2) $\widetilde{\underline{\textup{Gr}}}_{\mu}^{{\underline{\lambda}}_{N}}$$\underline{\textup{Gr}}_{\mu}^{{\underline{\lambda}}_{N}}$$\underline{\textup{Gr}}^{{\underline{\lambda}}_{N}}_{\mu,\mathbb{A}^{N}/\mathfrak{S}_{N}}$$\tilde{S}_{\mu}$$S_{\mu}\times_{\mathfrak{h}/W}\mathfrak{h}$$S_{\mu}$$\scriptstyle{\cong}\scriptstyle{\cong}\scriptstyle{\cong}$
@@ 3.3. Equivariant Slodowy slices
3.4. Zastava presentation of equivariant Slodowy slices in type AA
The cotangent bundle $T^{*}G$ for $G={{\textup{GL}}}_{N}$ has a realization as a Zastava space $Z^{\alpha_{0}^{*}}_{{{\textup{PGL}}}_{2N}}$ by [BP08], thus in particular a generalized slice in the affine Grassmannian of ${{\textup{PGL}}}_{2N}$. Further, by the Theorem 2.15 recalled above from [BFN19a], $T^{*}G$ is also realized as the Coulomb branch of a type $A_{2N}$ quiver, a linear quiver of length 2N12N-1. We now state this formally for future reference:
@@ 3.4. Zastava presentation of equivariant Slodowy slices in type AA
where $\bf{G}$} and $\bf{N}$} are as in the preceding Theorem 3.6, and in turn the definitions recalled in briefly in Section 2.6 above from [BFN18]. Since the resulting representation is trivial on the natural diagonal character $\mathbb{C}^{\times}\to{\bf{G}}={{\textup{GL}}}(V)=\prod_{i=1}^{N}{{\textup{GL}}}_{i}$, the above construction naturally descends to define a $\widetilde{\bf{G}}-$equivariant map $\mathcal{R}(\widetilde{\bf{G}},{\bf{N}})\xrightarrow{\tilde{\pi}_{F}}\mathrm{Gr}_{{{\textup{PGL}}}_{N}}$ where $\widetilde{\bf{G}}={\bf{G}}/\mathbb{C}^{\times}$. In summary, we obtain sheaves
@@
(3.4.2) $\operatorname{Spec}H^{\bullet}(\mathrm{Gr}_{{{\textup{PGL}}}_{N}},\mathcal{A}_{{{\textup{reg}}}})$ $=\mathcal{M}_{C}(\widetilde{\bf{G}},{\bf{N}})\xrightarrow{\cong}{{\textup{SL}}}_{N}\times S_{\textup{prin},\mathfrak{sl}_{N}}$ $\ ,$
@@
To begin, note that there is a manifest ${{\textup{GL}}}_{N}$ action on $\mathcal{M}_{C}({\bf{G}},{\bf{N}})$ as a factor of the group $L_{\textup{bal}}$ corresponding to the balanced subquiver (N,N1,N2,,1)(N,N-1,N-2,...,1). Moreover, this defines a trivial ${{\textup{GL}}}_{N}$-torsor, as we can directly calculate
@@
We also recall the following definition of the general genus zero Moore–Tachikawa varieties, with punctures labeled by dominant coweights $\mu^{1},...,\mu^{b}\in\Lambda^{+}_{{{\textup{PGL}}}_{N}}$ with $|\mu^{k}|=N$ for each k=1,,bk=1,...,b, or $\mu^{k}=0$ in which case the puncture is called ‘full’ and corresponds to those discussed in [BFN19b]:
@@ 4.1. Quantization in the algebraic setting
First, let us review some preliminaries on quantization. We largely follow the language of [BK04a, Los12]. For now, suppose that $Y$ is a smooth, symplectic, affine variety over $\mathbb{\mathbb{C}}$. By definition, ${\mathcal{O}}_{Y}$ is a sheaf of Poisson algebras over $Y$. Suppose also that $Y$ has an action of $\mathbb{G}_{m}$ such that the Poisson tensor is scaled with weight [^1] Usually one asks that $\mathbb{G}_{\hbar}$ scales the symplectic form and $\hbar$ with weight $2$, but one is then forced to change variables to $\sqrt{\hbar}$. We opt to avoid this by halving the weights. In characteristic zero, this amounts to little more than a matter of convention. 1-1, we use $\mathbb{G}_{\hbar}$ for this distinguished action. This group also also acts on $\mathbb{\mathbb{C}}{{[\![\hbar]\!]}}$ by scaling $\hbar$ with unit weight. Much of this discussion holds in a more general setting but in the spirit of brevity, we specialize to the setting we need. We use the term quantization for what we call a formal quantization in [BN25b].
Definition 4.1.
A quantization of the Poisson algebra ${\mathcal{O}}(Y)$ is a pair $(A,\varphi)$ of an associative $\mathbb{\mathbb{C}}{{[\![\hbar]\!]}}$ algebra $A$, flat over $\mathbb{\mathbb{C}}{{[\![\hbar]\!]}}$ and complete in the $\hbar$-adic topology, and an isomorphism of Poisson algebras $\varphi:A/\hbar A\rightarrow{\mathcal{O}}(Y)$.
A graded quantization is a pair $(A,\varphi)$ as above together with an action of $\mathbb{G}_{\hbar}$ on $A$ by algebra automorphisms, such that $\mathbb{G}_{\hbar}$ scales $\hbar$ with weight $1$ and $\varphi:A/\hbar A\rightarrow{\mathcal{O}}(Y)$ is $\mathbb{G}_{\hbar}$-equivariant.
Often, we will omit the isomorphism $\varphi$ in our notation when it is not relevant. These definitions have natural sheaf-theoretic analogues in quantizations of the structure sheaf ${\mathcal{O}}_{Y}$:
Definition 4.2.
A quantization of ${\mathcal{O}}_{Y}$ is a sheaf $\mathcal{A}$} on $Y$ of associative, flat, $\mathbb{\mathbb{C}}{{[\![\hbar]\!]}}$ algebras, complete in the $\hbar$-adic topology, and an isomorphism ${\mathcal{A}}/\hbar{\mathcal{A}}\rightarrow{\mathcal{O}}_{Y}$ of sheaves of Poisson algebras.
A graded quantization of ${\mathcal{O}}_{Y}$ is a pair $({\mathcal{A}},\varphi)$ as above, together with $\mathbb{G}_{\hbar}$-equivariant structure on the sheaf $\mathcal{A}$} such that $\mathbb{G}_{\hbar}$ scales $\hbar$ with weight $1$ and $\varphi$ is $\mathbb{G}_{\hbar}$-equivariant.
Note that the global sections of a graded quantization of ${\mathcal{O}}_{Y}$ are a graded quantization of ${\mathcal{O}}(Y)$.
@@
The moduli space of isomorphism classes of graded quantizations of ${\mathcal{O}}_{Y}$ is isomorphic to the second de Rham cohomology $H^{2}_{\rm dR}(Y)$.
This is the familiar result for quantizations of symplectic manifolds in the $C^{\infty}$ setting, see for example [dWL83, Fed94]. The result holds, in the algebraic context, for more general varieties but we only need the affine case.